# Emergence of Goal-Directed Behaviors via Active Inference with Self-Prior

### v1

@@ 3.1 Discrete Model
Here, we use uppercase letters (e.g., $P,\,Q$) to indicate discrete distributions, and denote categorical distributions by $\text{Cat}(\cdot)$. The observation $o_{t}$ and hidden state $s_{t}$ are one-hot column vectors with $N_{o}$ and $N_{s}$ rows, respectively, corresponding to the number of possible observations and hidden states. Boldface symbols such as $\mathbf{A}$ and $\mathbf{B}$ indicate the matrix parameters of the categorical distributions. $\mathbf{B}$ contains separate matrices for each available action $a$ (e.g., if actions are {LEFT, STOP, RIGHT}, then separate matrices $\mathbf{B}_{\text{LEFT}},\mathbf{B}_{\text{STOP}},\mathbf{B}_{\text{RIGHT}}$ exist). $\mathbf{A}$ is an $N_{o}\times N_{s}$ matrix, and each $\mathbf{B}_{a_{t-1}}$ is an $N_{s}\times N_{s}$ square matrix. Based on these definitions, the variational free energy $\mathcal{F}$ under a given action $a_{t-1}$ can be rewritten as:
$\mathcal{F}$ $=\mathbb{E}_{Q(s_{t})}[\log Q(s_{t})-\log P(o_{t},s_{t})]$ (5)
Accordingly, the parameter $\phi$ that minimizes free energy can be derived by taking the derivative of $\mathcal{F}$ with respect to $\phi$:
Here, $\sigma$ denotes the softmax function, which normalizes the elements of $\phi$ so that they sum to 1. In other words, the new posterior parameter $\phi$ can be computed from the previous action $a_{t-1}$ and the current observation $o_{t}$.
We assume that the likelihood distribution $\text{Cat}(\mathbf{A})$ and the prior distribution $\text{Cat}(\mathbf{B})$ have already been learned accurately. That is, the parameter $\mathbf{A}$ perfectly specifies the observation $o_{t}$ for each possible hidden state $s_{t}$, and $\mathbf{B}$ perfectly specifies the outcome of each possible action $a_{t}$ for each state. This simplifying assumption allows us to isolate the effect of changes in the self-prior on behavior, without introducing side effects from other parts of the generative model (in the continuous model discussed later, the learning of both likelihood and prior is included).
The distribution $\tilde{P}(o_{t})$, which serves as the setpoint for future observations, is given in the form of a categorical distribution $\text{Cat}(\mathbf{C})$, where $\mathbf{C}$ is a column vector with a row for each possible $o_{t}$. In this study, since we aim to examine behavior driven by the self-prior, all $o_{t}$ contain only information related to $o^{I}_{t}$. Therefore, the entire $\text{Cat}(\mathbf{C})$ is defined as the self-prior:
@@
The agent autonomously learns the self-prior by recording the frequency of previously experienced observations and storing the corresponding observation probabilities in the parameter $\mathbf{C}$. The initial value of $\mathbf{C}$ is set so that all observations have equal probability.
To determine the optimal action $a_{t}$, we compute the expected free energy associated with each possible action, as described above. This is given by:
For practical computation, the equation can be reformulated as follows:
Here, we use $Q(o_{t}\mid s_{t},a_{t-1})=P(o_{t}\mid s_{t})$ because the mapping from state to observation is assumed to be independent of action. Furthermore, since $\mathbf{A}$ and $\mathbf{B}$ are assumed to be known exactly in our setup, the action entropy term for parameter information gain is omitted in the discrete experiment (see (Parr et al.,, 2022) for a detailed explanation of above formulation).
In the discrete environment, we compute expected free energy starting from the current time $t$ to a future time $t+N$. The total expected free energy for all steps is then used to evaluate a policy $\pi=\{a_{t},a_{t+1},...,a_{t+N}\}$. The policy is computed as a categorical distribution over actions, from which actions are sampled as follows:
@@
3.2 Continuous Model
Prior research on extending the free energy principle from discrete to continuous domains often exploits the expressive power of deep neural networks. Indeed, several techniques originally proposed in deep reinforcement learning have been shown to closely approximate free energy optimization. For example, the Recurrent State-Space Model (RSSM) in PlaNet (Hafner et al.,, 2019) can be interpreted as optimizing variational free energy (Çatal et al.,, 2020), and policy gradient methods can be viewed as minimizing expected free energy (Millidge,, 2020). Following these precedents, we implement free energy minimization in a continuous environment using deep neural networks (Figure 2).
Whereas the discrete model fixed the parameters of the likelihood and prior distributions in variational free energy, the continuous model trains both distributions using deep neural networks, similar to the RSSM, with respect to the parameter $\phi$:
Prior: $p_{\phi}(h_{t}\mid s_{t-1},a_{t-1})$ (11)
Posterior: $q_{\phi}(s_{t}\mid h_{t},o_{t})$
Likelihood: $p_{\phi}(o_{t}\mid s_{t})$
The prior encodes the previous hidden state $s_{t-1}$ and action $a_{t-1}$ into a single vector via a linear layer, which is then fed into a GRU cell (Chung et al.,, 2014) to integrate inputs over time. The GRU’s hidden state is passed through a two-layer MLP that outputs the parameters of a multivariate Gaussian distribution, from which we sample a stochastic state. We concatenate this stochastic state with the GRU hidden state to produce the prior feature $h_{t}$. We use $h_{t}$ to distinguish the prior’s representation from the posterior variable $s_{t}$ (That is, the KL term in the variational free energy $\mathcal{F}$, originally $D_{\mathrm{KL}}[q(s_{t})\|p(s_{t})]$, is expressed as $D_{\mathrm{KL}}[q_{\phi}(s_{t}\mid h_{t},o_{t})\|p_{\phi}(h_{t}\mid s_{t-1},a_{t-1})]$).
@@ 3.2 Continuous Model
The posterior receives as input a concatenated vector of the prior feature $h_{t}$ and an embedded observation $o_{t}$. We use embedded observations because our experiment involves multimodal sensory data, and embedding allows us to integrate sensory inputs of different dimensionalities into a single representation. Since the embedding method may vary depending on the experiment, further details are provided in the Experiments section. The concatenated $h_{t}$ and $o_{t}$ are passed into a two-layer MLP in the posterior network, which outputs the parameters of a multivariate Gaussian distribution, and sample the posterior feature $s_{t}$ from this distribution.
The likelihood network transforms $s_{t}$ into $o_{t}$ through a linear layer, and the resulting embedded $o_{t}$ is then decoded into individual modality-specific observations. The parameters $\phi$ of the Prior, Posterior, and Likelihood networks are optimized via gradient descent to minimize the following free energy:
$\mathcal{F}$ $=\mathbb{E}_{q(s_{t})}[\log q(s_{t})-\log p(o_{t},s_{t})]$ (12)
$\therefore\operatorname*{\arg\min}_{\phi}\mathcal{F}$ $=\operatorname*{\arg\min}_{\phi}\big{[}D_{\mathrm{KL}}[q_{\phi}(s_{t}\mid h_{t},o_{t})\|p_{\phi}(h_{t}\mid s_{t-1},a_{t-1})]-\mathbb{E}_{q_{\phi}(s_{t}\mid h_{t},o_{t})}[\log p_{\phi}(o_{t}\mid s_{t})]\big{]}$
In the continuous model, as in the discrete case, we do not consider extrinsic preferences in order to focus on behavior driven by the self-prior. Thus, we set $\tilde{p}(o_{t})=\tilde{p}(o^{I}_{t})$. The self-prior is implemented using a normalizing-flows-based density estimator, specifically Neural Spline Flows (NSF) (Durkan et al.,, 2019). The agent autonomously learns the self-prior by maximizing the log-likelihood of the observation at each time step. For efficient training of neural networks, we sample observations from a replay buffer $\mathcal{D}$ that stores recently experienced trajectories:
@@
Unlike the discrete model, which plans actions by minimizing the sum of expected free energy up to a future time $t+N$, the continuous model cannot enumerate such possibilities. To address this, we adopt a learning approach inspired by the actor-critic method in reinforcement learning, similar to (Millidge,, 2020). Specifically, we train a separate policy network that predicts an appropriate action $a_{t}$ from the state $s_{t}$, and additionally train an expected utility network that estimates expected free energy over an infinite horizon beyond $t+N$:
@@
This approach is similar to Soft Actor-Critic (Haarnoja et al.,, 2018), where both the policy network and the utility network take the posterior feature $s_{t}$ as input, use a 3-layer MLP to output the parameters of a multivariate Gaussian distribution, and sample from this distribution. Accordingly, the policy and utility networks are trained as follows:
$G^{\lambda}_{t}$ $=\mathcal{G}(s_{t})+\gamma_{t}\begin{cases}(1-\lambda)g_{\psi}(s_{t+1})+\lambda G^{\lambda}_{t+1},&\text{if}\ t<H\\ g_{\psi}(s_{H}),&\text{if}\ t=H\end{cases}$
We use GAE($\lambda$) estimation to train the utility network (Schulman et al.,, 2015). For all other unspecified hyperparameters and training stabilization techniques throughout the continuous model (e.g., KL-divergence clipping, action entropy scaling), we follow (Mazzaglia et al.,, 2021).
@@ 4 Experiments and Results
In this section, we validate the emergence of goal-directed behavior driven by the self-prior, which is learned from multimodal sensory experiences, through computer simulations. Our experimental setup assumes a scenario in which an infant agent detects a sticker on its left arm, pays attention to the object, and eventually removes it with its right hand—similar to the study by (Bigelow,, 1986), where a silent toy was placed on the body of a blind infant to examine reaching behavior.
Our experiment is based on the following assumptions: (i) since an infant typically does not have stickers attached to its body, it acquires a self-prior that "there are no stickers on my body"; (ii) when a sticker is attached, the mismatch between sensory observations and the self-prior increases free energy; and (iii) to reduce free energy, the agent spontaneously generates reaching behavior aimed at the sticker.
Through these experiments, we (i) verify the process by which the self-prior, representing the agent’s own preference, is autonomously formed, (ii) confirm the emergence of reaching behavior within the free energy minimization framework incorporating the self-prior, and (iii) in the discrete environment, show that behavior can vary for the same stimulus depending on changes in the self-prior.
4.1 Discrete Simulation
In the discrete environment, the agent receives multimodal observations composed of discrete proprioceptive and tactile inputs (Figure 3). The agent’s right hand can occupy one of five positions, labeled $0,1,2,3,$ or $4$, representing its proprioceptive state. The left arm has three tactile sensors, located at positions $1,2,$ and $3$ in the right hand’s coordinate frame, and each sensor can be activated independently, yielding $82^{3}=$ 8 possible tactile states. Consequently, the observation variable $o_{t}$ can take $5\times 8=40$ distinct values and is represented as a 40-dimensional one-hot column vector (Table 1).
@@ 4.1 Discrete Simulation
At each time step, the agent selects one of three actions $a_{t}$: (i) remain stationary, (ii) move one step to the left, or (iii) move one step to the right. Thus, $a_{t}$ is represented as a 3-dimensional one-hot column vector. In the discrete experiment, the policy $\pi$ considers all $813^{4}=$ 81 possible action sequences over a 4-step future horizon, and each sequence is probabilistically chosen according to the sum of expected free energy as defined in Equation 10.
In the first phase of the discrete-environment experiment, we let the agent operate without constraints for a sufficiently long period (10,000 steps). In the case of real infants, such observations may correspond to a variety of everyday experiences that fulfill extrinsic preferences. However, for simplicity, we assumed a motor babbling condition in which the agent randomly selects its actions $a$. Initially, the self-prior was uniformly distributed across all possible observations, so it did not respond even when a sticker was present (Figure 5(a)).
Through motor babbling, the agent learned the pattern that “a single tactile signal occurs when the right hand is over the left arm,” and as a result, “no tactile signal occurs when the right hand is off the left arm” (Left of Figure 4). When a caregiver attaches a sticker to position 1 on the left arm, a tactile signal is generated at a location different from the current hand position (position 4). According to the newly learned self-prior, this is a low-probability situation, resulting in an increase in free energy (i.e., attention is drawn due to the mismatch with the model). The agent then begins planning actions to reduce expected free energy and concludes that touching the sticker aligns best with its self-prior. As a result, the agent shows goal-directed movement toward position 1 (Figure 5(b)). Similarly, when the caregiver moves the sticker to position 3, the agent detects this new signal as incongruent with its model and promptly reaches toward position 3.
@@
Figure 6: Agent’s behavior when a sticker has persistently been attached to position 3 ($t=30{,}000$). When the sticker is placed on other positions, the agent still reaches toward them; however, it no longer shows interest when the sticker is placed at position 3.
@@
4.2 Continuous Simulation
Up to this point, we have introduced the core idea that goal-directed behavior can emerge from the self-prior using a simple discrete setup. We now extend this to an environment with continuous-valued variables, inspired by the environment setup of (Marcel et al.,, 2022) (Figure 7).
[…]
To represent both tactile and proprioceptive modalities as a unified observation $o_{t}$ in the continuous model, we embed each modality into vectors of the same dimension. Specifically, we embed the tactile matrix using a four-layer Conv2D module and the proprioceptive vector using a two-layer MLP. The resulting embeddings are then combined via element-wise addition to yield the integrated observation embedding $o_{t}$. To decode from this embedding back to the original sensory modalities, we use a four-layer ConvTranspose2D for the tactile output and a two-layer MLP for the proprioceptive output.
[…]

### v2

@@ 3.1 Model for Discrete Environment
Here, uppercase letters P,QP,\,Q denote discrete distributions, and $\mathbf{A},\,\mathbf{B},\,\phi_{t}$ are parameter matrices and vectors of categorical distributions. $\mathbf{B}$ implements the transition model, with separate $\mathbf{B}_{a}$ matrices for each action $a$. Using this notation, minimizing variational free energy yields the optimal posterior distribution parameter $\phi_{t}$ for the current observation $o_{t}$ through the Softmax normalization function $\sigma$ as follows (derivation process see Appendix A.1.1):
$\phi_{t}$ $\approx\sigma(\log{(\mathbf{A}\cdot o_{t})}+\log{(\mathbf{B}_{a_{t-1}}\phi_{t-1})})$ (7)
We assume that the likelihood distribution $\text{Cat}(\mathbf{A})$ and the prior distribution $\text{Cat}(\mathbf{B})$ have already been learned accurately. That is, the parameter $\mathbf{A}$ perfectly knows which observation $o_{t}$ corresponds to each possible hidden state $s_{t}$, and $\mathbf{B}$ perfectly knows the outcome of each possible action $a_{t}$ for each state. This simplifying assumption is intended to exclude side effects that may arise from other parts of the generative model when investigating changes in behavior due to changes in the self-prior (the learning of both likelihood and prior distributions is included in the model for the continuous environment discussed later).
The distribution $\tilde{P}(o_{t})$ that serves as the setpoint for future observations is given in the form of a categorical distribution $\text{Cat}(\mathbf{C})$, where $\mathbf{C}$ is a column vector with a row for each possible $o_{t}$. According to the definition in Eq. (3), the preferred prior $\tilde{p}(o_{t})$ should decompose into an extrinsic component $\tilde{p}(o^{E}_{t})$ and an intrinsic component $\tilde{p}(o^{I}_{t})$, but in this study, we excluded the extrinsic component in order to investigate purely intrinsic motivational behavior driven by the self-prior. Therefore, all components of the observation vector $o_{t}$ correspond to intrinsic observations $o^{I}_{t}$, and the entire $\text{Cat}(\mathbf{C})$ is defined as the self-prior $\tilde{P}(o^{I}_{t})$:
@@
The agent autonomously learns the self-prior by recording the frequency of each observation experienced so far and storing the corresponding observation probabilities in the parameter $\mathbf{C}$. The initial value of $\mathbf{C}$ is set so that all observations have equal probability. To determine the optimal action $a_{t}$ based on this, we compute the expected free energy for each action (derivation process see Appendix A.1.2):
$\mathcal{G}\approx(\mathbf{B}_{a_{t}}\phi_{t})\cdot\mathcal{H}[\mathbf{A}]+D_{\mathrm{KL}}[\mathbf{A}\mathbf{B}_{a_{t}}\phi_{t}\|\mathbf{C}]$ (9)
To evaluate future policies, we propagate the current belief $\phi_{t}$ through the transition model $\mathbf{B}$ into future beliefs $\phi_{t+1:t+N}$ along "imagined time": $\phi_{t+\tau}=\mathbf{B}_{a_{t+\tau-1}}\phi_{t+\tau-1}$. The optimal policy is selected by summing expected free energies for each candidate policy $\pi=\{a_{t},...,a_{t+N}\}$. The policy is expressed as a categorical distribution, from which actions are sampled:
@@
3.2 Model for Continuous Environment
In prior research attempting to apply the free energy principle to (high-dimensional) continuous environments, many implementations use joint angles directly as internal states and generate actions through analytically derived backpropagation (e.g., (sancaktar_end_2020; priorelli_deep_2023)). While these approaches offer the advantage of analytical computation, they are limited in that the representation is constrained to joint angles, making it difficult to implicitly encode complex sensorimotor information in high-dimensional latent spaces and challenging to apply to complex tasks requiring long-term planning.
In contrast, approaches utilizing deep neural networks do not directly mimic biological details, but they provide a practical method to extend the core computational principles of active inference to high-dimensional problems (millidge_deep_2020). The focus of this study is to propose and demonstrate the conceptual mechanism of the self-prior, which is not bound to any specific neural implementation. Therefore, we follow the deep neural network based approach to enable representation of complex sensorimotor information in high-dimensional latent spaces and to perform future planning. Specifically, in this study, we follow the approach of prior research that interprets the Recurrent State-Space Model (RSSM) in PlaNet (hafner_planet_2019) as optimizing variational free energy (catal_learning_2020), and interprets policy gradient methods as minimizing expected free energy (millidge_deep_2020). The RSSM is essentially an extension of the Variational Autoencoder (VAE, (kingma2013auto)) architecture to sequential data, directly implementing the principle of variational free energy minimization.
Whereas the discrete model fixed the parameters of the likelihood and prior distributions in variational free energy, the continuous model follows the standard RSSM structure and trains the parameters $\phi$ using deep neural networks. The RSSM separates the deterministic latent state $h_{t}$ and the stochastic latent state $s_{t}$, enabling proper sequential decomposition:
Deterministic state: $h_{t}=f_{\phi}(h_{t-1},s_{t-1},a_{t-1})$ (11)
Stochastic prior: $p_{\phi}(\hat{s}_{t}\mid h_{t})$
Stochastic posterior: $q_{\phi}(s_{t}\mid h_{t},o_{t})$
Likelihood: $p_{\phi}(o_{t}\mid h_{t},s_{t})$
The deterministic state captures temporal dependencies through a GRU cell, while the stochastic prior and posterior each generate Gaussian distribution parameters from the deterministic states (and observations). This separation ensures proper variational decomposition at each time step (implementation details in Appendix A.2.2).
@@ 3.2 Model for Continuous Environment
The continuous model also investigates purely intrinsic motivational behavior, thus $\tilde{p}(o_{t})=\tilde{p}(o^{I}_{t})$. The self-prior is learned using Neural Spline Flows (NSF, (durkan_neural_2019)) to maximize observation log-likelihood. This is done by sampling observations $o$ from a replay buffer $\mathcal{D}$ that stores recently experienced trajectories:
@@
Since action candidates cannot be enumerated in continuous environments, we use policy gradient methods. Following (millidge_deep_2020), we train a policy network that predicts actions from states and a value network that estimates expected free energy over an infinite horizon:
@@
This is a similar approach to Soft Actor-Critic (haarnoja_sac_2018), and approximates expected utility through GAE($\lambda$) estimation (schulman_gae_2015). Training details and hyperparameters are provided in Appendix A.2.3 and (mazzaglia_contrastive_2021).
@@ 4 Experiments and Results
In this section, we validate the emergence of goal-directed behavior driven by the self-prior, which is learned from multimodal sensory experiences, through computer simulations. As briefly mentioned earlier, our experiment is based on the following assumptions: (i) since an infant typically does not have stickers attached to its body, it acquires a self-prior that "there are no stickers on my body"; (ii) when a sticker is attached, the mismatch between sensory observations and the self-prior increases free energy; and (iii) to reduce free energy, the agent spontaneously generates reaching behavior aimed at the sticker. Importantly, we did not preset any preferred states such as removing stickers or adopting specific postures; all goal-directed behaviors emerge from the self-prior autonomously formed from the sensations the agent experiences through motor babbling.
Through these experiments, we (i) verify the process by which the self-prior, representing the agent’s own preference, is autonomously formed, (ii) confirm the emergence of reaching behavior within the free energy minimization framework incorporating the self-prior, and (iii) in experiments in the discrete environment, additionally show that behavior can vary for the same stimulus depending on changes in the self-prior.
4.1 Simulation in Discrete Environment
In the discrete environment, the agent receives multimodal observations composed of discrete proprioceptive and tactile inputs (Figure 3). The agent’s right hand can occupy one of five positions, labeled $0,1,2,3,$ or $4$, representing its proprioceptive state. The left arm has three tactile sensors, located at positions $1,2,$ and $3$ in the right hand’s coordinate frame, and each sensor can be activated independently, yielding $2^{3}=8$ possible tactile states. Consequently, the observation variable $o_{t}$ can take $5\times 8=40$ distinct values and is represented as a 40-dimensional one-hot column vector (Table 2).
@@ 4.1 Simulation in Discrete Environment
At each time step, the agent can choose one of three actions $a_{t}$: (i) remain stationary, (ii) move one step to the left, (iii) move one step to the right. Thus, $a_{t}$ is represented as a 3-dimensional one-hot column vector. In the discrete environment experiment, there are $3^{4}=81$ possible policies $\pi$, which are sequences of actions up to 4 steps into the future, and are probabilistically determined by Equation 10 according to the sum of expected free energies for each.
At the beginning of the experiment, the self-prior was uniformly initialized across all possible observations. This makes the expected free energy equal for all policies, causing random policy selection. That is, no special response occurs even when a sticker is attached (Figure 5(a)).
Subsequently, we let the agent perform motor babbling without constraints for a sufficiently long period (10,000 steps). In the case of real infants, such experiences may correspond to a variety of everyday experiences that fulfill extrinsic preferences. However, for simplicity, we performed motor babbling in which the agent randomly selects its actions $a$. Through this process, the pattern that "a single tactile signal occurs when the right hand is over the left arm" is learned, and as a result, "no tactile signal occurs when the right hand is off the left arm" (Left of Figure 4).
At this point, when a caregiver attaches a sticker to position 1 on the left arm, a tactile signal occurs at a location different from the current hand position (position 4). According to the newly learned self-prior, this is a low-probability situation, resulting in an increase in free energy (i.e., attention is drawn due to the mismatch with the model). Accordingly, the agent begins planning actions to reduce expected free energy and concludes that touching the sticker aligns best with its self-prior. As a result, the agent shows goal-directed movement toward position 1 (Figure 5(b)). Similarly, when the caregiver moves the sticker to position 3, the agent detects this new signal as incongruent with its model and immediately reaches toward position 3.
@@
Figure 6: Agent’s behavior when a sticker has persistently been attached to position 3 (t=30,000t=30{,}000). When the sticker is placed on other positions, the agent still reaches toward them; however, it no longer shows interest when the sticker is placed at position 3.
@@
4.2 Simulation in Continuous Environment
Up to this point, we have introduced the core idea that goal-directed behavior can emerge from the self-prior using a simple discrete setup. We now extend this to an environment with continuous-valued variables, inspired by the environment setup of (marcel_learning_2022) (Figure 7).
Initially, the agent’s two arms lie on the $xy$-plane. The lengths of the left forearm, left arm, torso, right arm, and right forearm are $80$, $70$, $140$, $70$, and $80\,\text{mm}$, respectively. The left forearm features a $30\,\text{mm}$-wide two-dimensional tactile sensor array, placed laterally relative to the forearm itself. At the endpoint of the right forearm is a circular "right hand" of radius $6\,\text{mm}$, which can generate tactile input on the left forearm.
All joint angles are expressed in radians. The left elbow and left shoulder angles are fixed at $\pi/3$ and $2\pi/3$, while the right shoulder and right elbow angles have respective ranges of $0,\,2\pi/3]$ and $0,\,3\pi/4]$. Additionally, the right hand can move along a pseudo-$z$ (height) axis within the interval $0,\,20]\text{mm}$.
Similar to the discrete case, the agent perceives both proprioception and touch; however, these quantities are now continuous. The proprioceptive input is a three-dimensional real-valued vector $(\text{shoulder},\text{elbow},\text{hand})$, while tactile input is represented by an $80\times 30$ real-valued matrix normalized to [0, 1][0,\,1]. Whenever the right hand is above the left forearm at a height of $\leq 10\,\text{mm}$, the tactile sensor on the left forearm is activated with intensity inversely proportional to the hand’s height.
To represent both tactile and proprioceptive modalities as a unified observation $o_{t}$ in the model for the continuous environment, we convert each modality into an embedding of the same dimensionality and integrate them. The tactile matrix is transformed into an embedding using a CNN, and the proprioceptive vector is transformed into an embedding using a MLP. We combine the two embeddings via element-wise addition to yield the final integrated observation embedding $o_{t}$. To decode from the embedding back to the original sensations, we use inverse transformation modules (see Appendix A.2.1 for detailed hyperparameters).
A caregiver can attach a sticker of radius $4\,\text{mm}$ to the left forearm at a fixed height of $0\,\text{mm}$. As a simplified environmental rule to demonstrate the core self-prior mechanism, the agent can remove the sticker if the center of its right hand remains within the sum of the hand’s and sticker’s radii for $10$ consecutive time steps. This represents a simplified assumption wherein sustained contact between the hand and sticker leads to object removal, without modeling complex manipulation actions such as grasping or pushing.
At every time step, the right shoulder and right elbow angles can each rotate between 0.05-0.05 and $0.05\,\text{rad}$, and the hand’s height can shift by 0.5-0.5 to $0.5\,\text{mm}$. To keep the hand near the forearm, we impose an additional constraint that the hand’s center must remain within $15\,\text{mm}$ of the forearm’s surface. If a proposed movement would place the hand outside this region, the agent’s action is disregarded and its previous position is retained.